OpenStudy (abmon98):

Differentiate with respect to x 1/ln(x)

OpenStudy (anonymous):

\[f(x)=\frac{1}{ln(x)}\]\[f(x)=[ln(x)]^{-1}\]\[f'(x)=-\frac{1}{[ln(x)]^2}*\frac{d}{dx}ln(x)\]The answer is:\[f'(x)=-\frac{x}{[ln(x)]^2}\]

OpenStudy (anonymous):

final result is wrong @D3xt3R check it

OpenStudy (anonymous):

Ops, the correct answer is:\[f'(x)=-\frac{1}{x[ln(x)]^2}\]

OpenStudy (anonymous):

Thanks @shahzadjalbani

OpenStudy (anonymous):

You are welcome @D3xt3R

OpenStudy (abmon98):

Thank you @D3xt3R

OpenStudy (abmon98):

\[2x/\sqrt{x-1}\] \[2x*(x-1)^-1/2\] \[use product rule\] \[2*(x-1)^-1/2 + -1/2(x-1)^-3/2 *2x\] \[2/\sqrt{x-1)} -x/\sqrt{x-1}^3\] Is my working right

OpenStudy (abmon98):

How can i differentiate with respect to x \[2x \ln \sqrt{x+2}\]

geerky42 (geerky42):

apply product rule ( 2x and \(\ln\sqrt{x+2}\) ) for \(\ln\sqrt{x+2}\), you apply chain rule

OpenStudy (abmon98):

\[(2x)' *(\ln \sqrt{x+2})+(2x)*(\ln \sqrt{x+2})'\] \[2*(\ln \sqrt{x+2})+2x(1/x*1/2\sqrt{x+2})\]

geerky42 (geerky42):

actually: \((\ln\sqrt{x+2})'=(\ln\left[ (x+2)^{1/2}\right])'=\dfrac{1}{\sqrt{x+2}}\cdot\dfrac{1}{2\sqrt{x+2}}\)

geerky42 (geerky42):

so you just have \((\ln\sqrt{x+2})' = \dfrac{1}{2(x+2)}\)

geerky42 (geerky42):

does that make sense?

geerky42 (geerky42):

because chain rule states that \(\dfrac{d}{dx}f(~g(x)~) = f'(~g(x)~)~~g'(x)\)

OpenStudy (abmon98):

@geerky42 so you mean its the differentiate of the whole equation - the differentiation of sqrt of x+2

geerky42 (geerky42):

well, I applied chain rule by let f(x) be ln x and g(x) be \(\sqrt{x+2}\). so we have f'(g(x)) = \(\dfrac{1}{\sqrt{x+2}}\), because \((\ln x)' = \dfrac{1}{x}\), right? then we take derivative of g(x) and have it multiplied by f'(g(x)) and we have \(g'(x) = \dfrac{1}{2}(x+2)^{-1/2}=\dfrac{1}{2\sqrt{x+2}}\) So \(f'(g(x))\cdot g'(x) = \dfrac{1}{\sqrt{x+2}}\cdot\dfrac{1}{2\sqrt{x+2}} = \boxed{\dfrac{1}{2(x+2)}}\)

OpenStudy (abmon98):

Thank you so much :)

OpenStudy (abmon98):

Differentiate with respect to x 1/sin^2x

geerky42 (geerky42):

chain rule again: f(x) = 1/x, g(x) = sin^2 x find ( f(g(x)) )'